# Three-body molecular states composed of $D^{(*)}$ and two nucleons

### v1

@@ I Introduction
In recent years, heavy flavor hadrons have provided a crucial testing ground for exploring the nonperturbative regime of quantum chromodynamics (QCD). Experimentally, a number of exotic hadronic states containing charm or bottom quarks, have been observed, such as the $XYZ$ heavy-quarkonium-like states, $P_{c}$ pentaquark states and $T_{cc}$ tetraquark state. Representative experimental measurements can be found in Refs. [23, 6, 50, 1, 3, 5, 4]. For comprehensive reviews on this rapidly developing field, we refer the reader to Refs. [66, 69, 15, 16, 42, 20, 30, 46, 11, 36, 59, 49, 17, 58, 19, 71] . In this context, systems composed of open heavy-flavor mesons, such as $D$ and $D^{*}$, and nucleons have attracted considerable attention, as they provide a unique setting in which heavy-quark spin symmetry (HQSS) and light-quark chiral dynamics interplay, potentially supporting unconventional bound or resonant states.
@@
In contrast to the well-established $NN$ interaction, the study of heavy-meson–nucleon systems, such as $DN$ and $D^{*}N$, is comparatively recent but has progressed rapidly in the past two decades. On the theoretical side, effective interactions constrained by heavy-quark symmetry and chiral symmetry for light quarks have been systematically developed. Early investigations based on chiral unitary and coupled-channel approaches demonstrated that near-threshold structures can be dynamically generated in the $DN$ and $D^{*}N$ systems [51, 33]. Subsequent studies employing one-boson-exchange models and effective Lagrangians further emphasized the role of tensor forces and channel coupling in providing attraction in selected quantum-number channels [39, 37].
@@
Combining the well-established $NN$ potentials with the experimentally constrained $DN$ interactions naturally motivates the study of few-body systems containing a heavy meson and multiple nucleons. The three-body system $DNN$, with $D$ a charmed meson and $N$ a nucleon, provides the simplest nontrivial environment to explore how heavy-quark symmetry, realistic nuclear correlations, and meson-baryon dynamics cooperate in forming heavy-flavor few-body states [34]. Earlier variational and Faddeev-type studies (including fixed-center approximations) pointed out that the $DNN$ system may support a quasibound structure with moderate binding and a narrow width in selected channels [25]. In the present work we go beyond this qualitative expectation by employing an energy-independent $D^{(*)}N$ interaction constrained by heavy-quark spin symmetry together with a realistic $NN$ potential, and by solving the three-body Schrödinger equation using the Gaussian Expansion Method (GEM) [40]. Possible resonant structures are further analyzed within the Complex Scaling Method (CSM) [9, 12]. The combined application of GEM and CSM has proven highly effective in studying few-body resonant states across various scales, such as in few-hadron systems [88, 72, 21], few-lepton systems [52, 73], and multiquark states [22, 53, 57, 77, 79, 78].
It is also instructive to recall that few-body systems composed of a meson and two nucleons have been extensively investigated in the strange sector. In particular, the $\bar{K}NN$ ($K^{-}pp$) system has been studied using variational and Faddeev-type approaches, where the strong $\bar{K}N$ interaction can generate quasibound or deeply bound three-body states depending on the treatment of coupled channels and short-range dynamics [10, 64, 26, 31, 67]. Although the predicted binding energies and decay widths remain model-dependent, these studies illustrate a general mechanism: realistic $NN$ correlations combined with meson–nucleon attraction could produce nontrivial three-body structures and pronounced spatial rearrangement. Experimental indications of compact $\bar{K}NN$ configurations have also been reported in kaon-induced and proton-proton reactions [8, 82]. This motivates analogous explorations in the charm sector, where the heavier meson mass reduces kinetic energy and heavy-quark spin symmetry enforces strong channel coupling between $DN$ and $D^{*}N$, potentially favoring the formation of bound states with distinctive spin dependence.
@@ III.1 $D^{(*)}N$ system: binding energies and spatial structure
Table 2: Comparison of the binding energies (in MeV) for the $S$-wave DNDN and $D^{*}N$ systems with isospin $I$ and total spin $S$, obtained in different approaches. The results of this work correspond to the cutoff $\Lambda=1.4~\mathrm{GeV}$ with $R_{\lambda}=R_{\beta}=R_{s}=1$.The RMS radii are given in units of fm.
System $I(J^{P})$ ChEFT [70] ChQM [27] QDCSM [86] This work $R_{rms}$
DNDN $0(1/2^{-})$ $11.1$ 1.70 $2.22$ $2.79$
$1(1/2^{-})$
$D^{*}N$ $0(1/2^{-})$ $1.5$
$1(1/2^{-})$ 0.48 $0.57$ $4.47$
$0(3/2^{-})$ $6.7$ 8.02 3.4 $9.75$ $1.62$
$1(3/2^{-})$
Table 2 presents a systematic comparison of the binding energies of the $S$-wave DNDN and $D^{*}N$ systems with different isospin $I$ and total spin $S$ obtained in various theoretical approaches. The results of the present work are shown together with those from chiral effective field theory (ChEFT) [70], chiral quark models (ChQM) [27], and the quark delocalization color screening model (QDCSM) [86], providing a common basis for comparing the near-threshold behavior predicted by different theoretical approaches. We note that the qualitative features discussed below are not sensitive to the specific choice of the rescaling factors; for simplicity and clarity of presentation, all rescaling factors are set to unity in Table 2.
@@
In contrast, the isovector $DN$ channel with $I(J^{P})=1(\tfrac{1}{2}^{-})$ remains unbound in the present calculation. This finding is consistent with the conclusions of ChEFT [70] and QDCSM-based studies [86], although some quark-model calculations predict attraction in this channel for specific parameter choices [27, 84]. The absence of a bound-state solution in this channel within the present framework further disfavors a molecular interpretation of the isovector $DN$ system.
@@
The $D^{*}N$ interaction has also been investigated within the one-boson-exchange (OBE) framework [39]. While the present results are qualitatively consistent with these OBE studies, differences in the cutoff dependence are observed. In particular, bound-state solutions in certain $D^{*}N$ channels appear only when the cutoff parameter exceeds a moderate value, indicating that the detailed quantitative predictions remain sensitive to the regularization scheme. Such differences suggest that the discrepancies among various OBE-based calculations originate primarily from the treatment of short-range regularization.
To further illustrate the cutoff dependence of the bound-state solutions obtained in the present work, we summarize in Tables 3 and 4 the binding energies and RMS radii of the $DN$ and $D^{*}N$ systems for $\Lambda=1.35-1.45~\mathrm{GeV}$. As shown in the table, the isoscalar $DN$ channel with $J^{P}=\tfrac{1}{2}^{-}$ develops a shallow bound state as the cutoff exceeds $\Lambda\simeq 1.35~\mathrm{GeV}$, with the binding energy increasing and the RMS radius decreasing as $\Lambda$ increases. A similar but more pronounced behavior is observed for the isoscalar $D^{*}N$ channel with $J=\tfrac{3}{2}$, which exhibits a stable bound-state solution already at $\Lambda=1.35~\mathrm{GeV}$ and remains relatively compact over the considered cutoff range. In contrast, the $I=1$ channels either remain unbound or show only extremely weak attraction accompanied by large RMS radii, indicating near-threshold or loosely correlated configurations rather than genuine molecular bound states.
@@
III.2 $D^{(*)}NN$ system: binding energies and spatial structure
[…]
The resulting binding energies and RMS radii of the $D^{(*)}NN$ systems are summarized in Figs. 3 and 4, respectively. We first discuss the $DNN$ system. For $\Lambda=1.1~\mathrm{GeV}$, the three-body binding energy decreases from $47.76~\mathrm{MeV}$ at a $Z_{c}$ pole of $5~\mathrm{MeV}$ to $13.26~\mathrm{MeV}$ at $35~\mathrm{MeV}$. A similar monotonic behavior is observed for $\Lambda=1.2$ and $1.3~\mathrm{GeV}$, where the binding energies span the ranges $59.71$-$23.17~\mathrm{MeV}$ and $63.62$-$31.03~\mathrm{MeV}$, respectively. These values are substantially larger than the binding energy of the two-body $NN$ system alone, $B_{NN}=2.2~\mathrm{MeV}$.
[…]
To further clarify the nature of the bound states obtained in the present calculation, we additionally examine the pole trajectories of the three-body $D^{(*)}NN$ systems in the complex energy plane. The results are shown in Fig. 5, where the real and imaginary parts of the three-body energies are displayed for the $DNN$ and $D^{*}NN$ systems with $J^{P}=0^{-}$, $1^{-}$, and $2^{-}$. These calculations are performed within the rescaling scheme with $R_{s}=R_{\lambda}=R_{\beta}=1$ and a fixed cutoff $\Lambda=1.4~\mathrm{GeV}$.
A salient feature of Fig. 5 is that all poles associated with the $DNN$ and $D^{*}NN$ systems move strictly along the negative real-energy axis without developing a sizable imaginary part. This behavior demonstrates that the states obtained in the present work correspond to genuine bound states rather than resonant or virtual states. In particular, no pole is observed to cross into the lower-half complex plane away from the real axis, which would be characteristic of a three-body resonance. The absence of such behavior provides strong evidence that the deeply bound states found in Figs. 3 and 4 are not artifacts of threshold effects or numerical continuation, but represent true bound solutions of the three-body problem.
IV Conclusion
Acknowledgements.
This work is supported by the Special Funds for Theoretical Physics under the National Natural Science Foundation of China (Grant No. 12547105), the National Natural Science Foundation of China (Grant No. 12575153) and the Start-up Funds of Southeast University (Grant No. 4007022506)

### v2

@@ I Introduction
In recent years, heavy flavor hadrons have provided a crucial testing ground for exploring the nonperturbative regime of quantum chromodynamics (QCD). Experimentally, a number of exotic hadronic states containing charm or bottom quarks, have been observed, such as the $XYZ$ heavy-quarkonium-like states, $P_{c}$ pentaquark states and $T_{cc}$ tetraquark state. Representative experimental measurements can be found in Refs. [23, 6, 52, 1, 3, 5, 4]. For comprehensive reviews on this rapidly developing field, we refer the reader to Refs. [70, 72, 15, 16, 43, 20, 30, 47, 11, 37, 62, 51, 17, 61, 19, 49, 74] . In this context, systems composed of open heavy-flavor mesons, such as $D$ and $D^{*}$, and nucleons have attracted considerable attention, as they provide a unique setting in which heavy-quark spin symmetry (HQSS) and light-quark chiral dynamics interplay, potentially supporting unconventional bound or resonant states.
@@
In contrast to the well-established $NN$ interaction, the study of heavy-meson–nucleon systems, such as $DN$ and $D^{*}N$, is comparatively recent but has progressed rapidly in the past two decades. On the theoretical side, effective interactions constrained by heavy-quark symmetry and chiral symmetry for light quarks have been systematically developed. Early investigations based on chiral unitary and coupled-channel approaches demonstrated that near-threshold structures can be dynamically generated in the $DN$ and $D^{*}N$ systems [54, 34]. Subsequent studies employing one-boson-exchange models and effective Lagrangians further emphasized the role of tensor forces and channel coupling in providing attraction in selected quantum-number channels [40, 38]. Recently, the $ND$ scattering lengths was extracted from the $\Lambda_{b}\to\pi^{-}pD^{0}$ decays by making use of the cusp at the $nD^{+}$ threshold.[66]
@@
Combining the well-established $NN$ potentials with the experimentally constrained $DN$ interactions naturally motivates the study of few-body systems containing a heavy meson and multiple nucleons. The three-body system $DNN$, with $D$ a charmed meson and $N$ a nucleon, provides the simplest nontrivial environment to explore how heavy-quark symmetry, realistic nuclear correlations, and meson-baryon dynamics cooperate in forming heavy-flavor few-body states [35].Beyond charm–nuclear systems, purely heavy-meson three-body molecular configurations have also been investigated, where heavy-quark symmetry and short-range dynamics can generate nontrivial bound structures in systems composed entirely of heavy mesons [53, 33, 80]. Earlier variational and Faddeev-type studies (including fixed-center approximations) pointed out that the $DNN$ system may support a quasibound structure with moderate binding and a narrow width in selected channels [25]. In the present work we go beyond this qualitative expectation by employing an energy-independent $D^{(*)}N$ interaction constrained by heavy-quark spin symmetry together with a realistic $NN$ potential, and by solving the three-body Schrödinger equation using the Gaussian Expansion Method (GEM) [41]. Possible resonant structures are further analyzed within the Complex Scaling Method (CSM) [9, 12]. The combined application of GEM and CSM has proven highly effective in studying few-body resonant states across various scales, such as in few-hadron systems [92, 75, 21], few-lepton systems [55, 76], and multiquark states [22, 56, 60, 81, 83, 82].
It is also instructive to recall that few-body systems composed of a meson and two nucleons have been extensively investigated in the strange sector. In particular, the $\bar{K}NN$ ($K^{-}pp$) system has been studied using variational and Faddeev-type approaches, where the strong $\bar{K}N$ interaction can generate quasibound or deeply bound three-body states depending on the treatment of coupled channels and short-range dynamics [31, 10, 68, 26, Topolnicki:2015zya,Zhang:2021hcl]. Although the predicted binding energies and decay widths remain model-dependent, these studies illustrate a general mechanism: realistic $NN$ correlations combined with meson–nucleon attraction could produce nontrivial three-body structures and pronounced spatial rearrangement. Experimental indications of compact $\bar{K}NN$ configurations have also been reported in kaon-induced and proton-proton reactions [8, 86]. This motivates analogous explorations in the charm sector, where the heavier meson mass reduces kinetic energy and heavy-quark spin symmetry enforces strong channel coupling between $DN$ and $D^{*}N$, potentially favoring the formation of bound states with distinctive spin dependence.
@@ III Numerical Results and Analysis
Table 2: Comparison of the binding energies (in MeV) for the $S$-wave DNDN and $D^{*}N$ systems with isospin $I$ and total spin $S$, obtained in different approaches. The results of this work correspond to the cutoff $\Lambda=1.2~\mathrm{GeV}$. The RMS radii are given in units of fm. Here “ub” denotes that the corresponding channel is unbound. In this work, numerical results spanning a range mainly due to couplings constrained by the $Z_{c}(3900)$ virtual-state pole variation.
System $I(J^{P})$ ChEFT [73] ChQM [27] QDCSM [90] This work $R_{rms}$
DNDN $0(1/2^{-})$ $11.1$ 1.70 0.350.354.544.54 2.232.235.595.59
$1(1/2^{-})$ ub – $2.72$ ub – $5.79$
$D^{*}N$ $0(1/2^{-})$ $1.5$
$1(1/2^{-})$ 0.48 0.820.8225.8425.84 1.061.064.364.36
$0(3/2^{-})$ $6.7$ 8.02 3.4 8.358.3543.2343.23 0.880.881.701.70
$1(3/2^{-})$ ub – $0.33$ ub – $5.76$
@@ III.1 $D^{(*)}N$ system: binding energies and spatial structure
Table 2 presents a systematic comparison of the binding energies of the $S$-wave DNDN and $D^{*}N$ systems with different isospin $I$ and total spin $S$ obtained in various theoretical approaches. The results of the present work are shown together with those from chiral effective field theory (ChEFT) [73], chiral quark models (ChQM) [27], and the quark delocalization color screening model (QDCSM) [90], providing a common basis for comparing the near-threshold behavior predicted by different theoretical approaches.
@@
In contrast, the isovector $DN$ channel with $I(J^{P})=1(\tfrac{1}{2}^{-})$ remains unbound in the present calculation. This finding is consistent with the conclusions of ChEFT [73] and QDCSM-based studies [90], although some quark-model calculations predict attraction in this channel for specific parameter choices [27, 88]. The absence of a bound-state solution in this channel within the present framework further disfavors a molecular state of the isovector $DN$ system.
However, whether the isovector $DN$ channel with $I(J^{P})=1(\tfrac{1}{2}^{-})$ remains unbound is inconclusive in the present calculation. When adopting different $DD\sigma$ couplings constrained by the variation of the $Z_{c}(3900)$ virtual-state pole, the results shift between bound and unbound states. Nevertheless, even in the unbound scenarios, the interaction extracted in this work for this channel is attractive. In the literature, no bound state is found in this channel in ChEFT [73] or QDCSM-based studies [90]. Notably, however, some quark-model calculations predict the existence of a bound state [27, 88], as also indicated in Refs. [66, 38].
@@
The $D^{*}N$ interaction has also been investigated within the one-boson-exchange (OBE) framework [40, 38]. While the present results are qualitatively consistent with these OBE studies, differences in the cutoff dependence are observed. In particular, bound-state solutions in certain $D^{*}N$ channels appear only when the cutoff parameter exceeds a moderate value, indicating that the detailed quantitative predictions remain sensitive to the regularization scheme. Such differences suggest that the discrepancies among various OBE-based calculations originate primarily from the treatment of short-range regularization.
To further illustrate the regulator and input dependence of the two-body sector, we show in Figs. 3 and 4 the binding energies and RMS radii of the $S$-wave DNDN and $D^{*}N$ systems as functions of the $Z_{c}(3900)$ virtual-state pole position, for three representative cutoffs $\Lambda=1.1$, $1.2$, and $1.3~\mathrm{GeV}$. As the $Z_{c}(3900)$ pole is moved away from the threshold (from 5-5 to $35~\mathrm{MeV}$), the $DD\sigma$ couplings become weaker, leading to systematically $D^{(*)}N$ reduced binding energies and increased spatial extents in all channels that exhibit attraction. This trend provides a transparent diagnostic of how the two-body inputs propagate into the three-body calculations. For the isoscalar $DN$ channel with $I(J^{P})=0(\tfrac{1}{2}^{-})$, the interaction remains only weakly attractive in the explored parameter region. A shallow bound state can appear for moderate-to-large cutoffs and for relatively shallow $Z_{c}$ poles, while it quickly dissolves when the couplings are weakened by shifting the $Z_{c}$ pole to more negative energies. Correspondingly, the RMS radius increases rapidly in the near-threshold region, indicating a spatially extended configuration characteristic of a weakly bound state. In contrast, the isovector $DN$ channel $I(J^{P})=1(\tfrac{1}{2}^{-})$ does not support a bound-state solution for the baseline parameters in the entire range shown in the figures, although the overall trend indicates that the interaction is still mildly attractive, consistent with the discussion above. A more robust pattern is observed in the $D^{*}N$ sector. The isoscalar channel with $I(J^{P})=0(\tfrac{3}{2}^{-})$ exhibits the strongest and most stable binding among the channels considered: it remains bound throughout a broad range of $Z_{c}$ pole positions and cutoffs, and its RMS radius stays comparatively small, pointing to a compact molecular configuration. By contrast, the isovector $D^{*}N$ channel with $I(J^{P})=1(\tfrac{1}{2}^{-})$ is much more sensitive to the short-range regularization and the fitted couplings: it typically appears as a very shallow bound state only for sufficiently strong couplings (i.e., for a shallow $Z_{c}$ pole) and/or larger $\Lambda$, and is accompanied by large RMS radii, suggesting a near-threshold and spatially diffuse structure. No bound-state solution is found for the remaining $I=1$ $D^{*}N$ channels in the parameter region displayed. These two-body trends provide the baseline for the three-body analysis below. In particular, the coexistence of a robustly bound $D^{*}N$ isoscalar channel and several near-threshold (or unbound but attractive) channels implies that three-body correlations can substantially enhance binding and induce spatial compression even when individual two-body subsystems are only weakly bound.
@@
Figure 3: Binding energies of the $S$-wave DNDN and $D^{*}N$ systems as functions of the $Z_{c}(3900)$ virtual-state pole position, for three representative cutoff values $\Lambda=1.1$, $1.2$, and $1.3~\mathrm{GeV}$. Each column corresponds to a fixed cutoff. The overall trend illustrates how the two-body binding energies evolve with the input pole position used to constrain the interaction.
Figure 4: Root-mean-square (RMS) radii of the $S$-wave DNDN and $D^{*}N$ systems as functions of the $Z_{c}(3900)$ virtual-state pole position, for the same cutoff values as in Fig. 3. The spatial sizes increase as the binding energies decrease, indicating near-threshold and spatially extended configurations.
III.2 $D^{(*)}NN$ system: binding energies and spatial structure
[…]
The resulting binding energies and RMS radii of the $D^{(*)}NN$ systems are summarized in Figs. 5 and 6, respectively. For a transparent comparison with the corresponding three-body states, we display in each partial-wave channel the most deeply bound two-body state among the $D^{(*)}N$ subsystems as a reference. We first discuss the $DNN$ system. For $\Lambda=1.1~\mathrm{GeV}$, the three-body binding energy decreases from $47.76~\mathrm{MeV}$ at a $Z_{c}$ pole of $5~\mathrm{MeV}$ to $13.26~\mathrm{MeV}$ at $35~\mathrm{MeV}$. A similar monotonic behavior is observed for $\Lambda=1.2$ and $1.3~\mathrm{GeV}$, where the binding energies span the ranges $59.71$-$23.17~\mathrm{MeV}$ and $63.62$-$31.03~\mathrm{MeV}$, respectively. These values are substantially larger than the binding energy of the two-body $NN$ system alone, $B_{NN}=2.2~\mathrm{MeV}$.
[…]
To further clarify the nature of the bound states obtained in the present calculation, we additionally examine the pole trajectories of the three-body $D^{(*)}NN$ systems in the complex energy plane. The results are shown in Fig. 7, where the real and imaginary parts of the three-body energies are displayed for the $DNN$ and $D^{*}NN$ systems with $J^{P}=0^{-}$, $1^{-}$, and $2^{-}$. We note that the qualitative features discussed below are not sensitive to the specific choice of the rescaling factors. For simplicity and clarity of presentation, we set $R_{s}=R_{\lambda}=R_{\beta}=1$ and adopt a fixed cutoff $\Lambda=1.4~\mathrm{GeV}$ in Fig. 7.The corresponding numerical results are summarized in Tables. 3 and 4 in the Appendix.
A salient feature of Fig. 7 is that all poles associated with the $DNN$ and $D^{*}NN$ systems move strictly along the negative real-energy axis without developing a sizable imaginary part. This behavior demonstrates that the states obtained in the present work correspond to genuine bound states rather than resonant or virtual states. In particular, no pole is observed to cross into the lower-half complex plane away from the real axis, which would be characteristic of a three-body resonance. The absence of such behavior provides strong evidence that the deeply bound states found in Figs. 5 and 6 are not artifacts of threshold effects or numerical continuation, but represent true bound solutions of the three-body problem.
IV Conclusion
Acknowledgements.
We are grateful to Feng-Kun Guo, Ming-Zhu Liu, Si-Qiang Luo, Tian-Wei Wu, and Hai-Peng Xie for useful discussions. This work is supported by the Special Funds for Theoretical Physics under the National Natural Science Foundation of China (Grant No. 12547105), the National Natural Science Foundation of China (Grant No. 12575153) and the Start-up Funds of Southeast University (Grant No. 4007022506).
Appendix A Numerical details of the $D^{(*)}NN$ and $D^{(*)}N$ system